The acceleration of a bus is given by 𝑎𝑥(𝑡) = (1.2 m/s3) t.

(a) If the bus’ velocity at time t=1.0 s is

5.0 m/s, what is its velocity at time t=2.0 s?

(b) If the bus’ position at time t=1.0 s is 6.0 m, what is

its position at time t=2.0 s?


1
Expert's answer
2021-03-08T08:27:25-0500

(a) Acceleration is the derivative of the velocity with respect to time:


"a=\dfrac{dv}{dt}."

Then, we can find the velocity of the bus by integrating "a_x(t)" over time:


"\displaystyle\intop_{v_0}^v dv=\displaystyle\intop_{0}^t adt,""v-v_0=1.2\displaystyle\intop_{0}^t tdt,""v=v_0+0.6t^2."

Let's first find "v_0" at t=1.0 s:


"v_0=v-0.6t^2=5\dfrac{m}{s}-0.6\cdot(1)^2=4.4\ \dfrac{m}{s}."

Finally, we can find velocity of the bus at t=2.0 s:


"v(t=2\ s)=4.4\ \dfrac{m}{s}+0.6\cdot(2\ s)^2=6.8\ \dfrac{m}{s}."

(b) Velocity is the derivative of the position with respect to time:


"v=\dfrac{dx}{dt}."

Then, we can find the position of the bus by integrating "v_x(t)" over time:


"\displaystyle\intop_{x_0}^x dx=\displaystyle\intop_{0}^t vdt,""x-x_0=\displaystyle\intop_{0}^t v_0dt+\displaystyle\intop_{0}^t 0.6t^2dt,""x=x_0+v_0t+0.2t^3."

Let's first find the inititial positon of the bus at t=1.0 s:


"x_0=6.0\ m-4.4\ \dfrac{m}{s}\cdot1\ s-0.2\cdot(1\ s)^3=1.4\ m."

Finally, we can find position of the bus at t=2.0 s:


"x(t=2\ s)=1.4\ m+4.4\ \dfrac{m}{s}\cdot2\ s+0.2\cdot(2\ s)^3=11.8\ m."